OpenStudy (anonymous):

Katalin drove 210 miles on her vacation. She drove an average of 1.4 times faster on the second 105 miles of her trip than she did on the first 105 miles of her trip. Which expression represents the time she spent driving? Let x = her speed on the first half of the trip.PLEASE HELP!

OpenStudy (anonymous):

T = D/S (time = distance divided by speed) T = 105/x + 105/(x + 1.4) I am not sure if this is correct....do you have answer choices ?

OpenStudy (anonymous):

@Opcode ....did I set this up correct ?

OpenStudy (anonymous):

@Luis_Rivera ....did I set this up correctly ?

OpenStudy (anonymous):

@Mertsj ....did I set this up correctly ?

OpenStudy (anonymous):

@Preetha .....did I set it up right ?

OpenStudy (anonymous):

@jim_thompson5910 .....@satellite73 ....did I set this up correctly ?

OpenStudy (anonymous):

@satellite73 .....help

jimthompson5910 (jim_thompson5910):

Let x = her speed on the first half of the trip (given) "She drove an average of 1.4 times faster on the second 105 miles of her trip than she did on the first 105 miles of her trip" means that Speed on second half of trip = 1.4x

OpenStudy (anonymous):

oh...I was adding it , not multiplying....oops

jimthompson5910 (jim_thompson5910):

First half d = rt 105 = x*t1 t1 = 105/x ------------------------------------------------------- Second half d = rt 105 = 1.4x*t2 t2 = 105/(1.4x) t2 = 75/x

jimthompson5910 (jim_thompson5910):

The two times t1 and t2 add to t1 + t2 105/x + 75/x (105+75)/x 180/x So the total time spent on the trip is 180/x hours

OpenStudy (anonymous):

you are really good.....thanks so much

jimthompson5910 (jim_thompson5910):

yw